(a) There is an urn containing six balls, three of which are red. You make two random draws from the urn without replacement. What is the probability that both of the balls you draw are red?
Answer #1

a) Sampling is being done without replacement. Hence,

P(Both red) = (3/6)*(2/5) = 1/5